# SCILLA: SurfaCe Implicit Learning for Large urban Area, a volumetric hybrid solution

### v3

@@ 3.3 SCILLA’s volume rendering
The strategy of associating a sample to either a volumetric or surface-based representation is discussed in the next section.
@@
SDF progressive initialization: Volumetric to surface representation. Starting from a pure volumetric density training ($\alpha_{i}=\alpha_{i}^{v}$), we introduce a progressive initialization of the signed distance field, that transitions from a volumetric to an SDF representation. Our approach is based on the premise that estimating the density $\sigma_{i}$ at a specific location requires less complex geometric understanding in contrast to estimating the signed distance to the closest point to the surface (i.e. $f(x_{i})$). 3D geometry represented as density field contains less information than SDF field: the former contains occupancy information (which can be considered as binary) whereas the latter represents in addition continuous distances to the surface. Therefore, learning a density field is simpler compared to the SDF field (see Fig. 6 of our ablation section). By initially leveraging a pure volumetric density representation during training, we shorten the convergence time of our density estimator.
@@
  • Volumetric stage: all the samples along the ray express volumetric densities, and their alpha values are computed with $\sigma_{i}$.
  • Hybrid stage: we gradually assign samples along the ray to train our signed distance field. Within the same ray, a portion of sample values is represented with volumetric alpha $\alpha_{i}^{v}$, while the remainder is computed with $\alpha_{i}^{f}$.
@@
Hybrid stage regularization. Composing alpha values obtained from our two different geometric representations is not straightforward as the learning dynamic of the density and the SDF are not the same. As mentioned earlier, learning a volumetric representation of the scene based only on spatial density is easier and faster to learn compared to learning an arbitrary signed distance field of a complex outdoor scene. To facilitate the convergence of our representation, we introduce two regularization methods: probability-based sampling attribution and SDF-gradient normalization.
Probability-based samples attribution. To assign the alpha value based on density or SDF to a sample, we initially start with a random selection for each ray. We noticed that this association strategy is not suitable: the signed distance field could not be properly initialized during the hybrid stage. To solve this problem, we introduce a probability-based sample association. We assign $\alpha_{i}^{f}$ to samples with the highest predicted density to initialize the signed distance field. It facilitates the learning of the signed distance field by limiting the range of predicted values, as all the points are sampled close to a surface.
SDF-gradient normalization. During the hybrid stage, where $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are jointly composed, we observed that the surface representation tends to predict large SDF gradient values to align $\alpha_{i}^{f}$ with $\alpha_{i}^{v}$ (we refer to our supplementary materials for further details). To address this, we normalize the SDF gradient in the $\alpha_{i}^{f}$ computation to prevent it from compensating for differences in the alpha distribution.
@@ 3.4 Optimization
According to NeuS formulation of the alpha compositing equation (Eq. 3), the network converges properly when the standard deviation $1/s$ of $\phi_{s}$ approaches zero. In complex driving scenarios and very large scenes, the parameter $s$ does not always reach a satisfactory level of convergence. Unlike StreetSurf which manually fixes the parameter $s$ to increment linearly, we introduce a regularization term on $s$: $\mathcal{L}_{\text{s}}=1/(s+\epsilon)$.
KITTI-360 [20] Pandaset [41] Waymo [32] nuScenes [3]
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Mean cumulative delta error for both SCILLA and StreetSurf computed across the four sequences from each dataset as detailed in Tab. 1. Curves indicate the percentage of ground-truth points having an error distance to the closest predicted mesh triangle which is lower than a given value. Light contours represent the standard deviation for each method.
@@
StreetSurf [11] 0.19 0.50 0.15 0.71 0.16 0.67 0.15 0.66 2.35 0.10 0.48 0.4 0.40 0.30 0.28 0.31
SCILLA (ours) 0.18 0.56 0.15 0.71 0.15 0.66 0.16 0.72 0.51 0.23 0.29 0.46 0.36 0.29 0.35 0.34
COLMAP [30] 0.15 0.70 0.12 0.81 0.12 0.78 0.15 0.74 0.71 0.41 1.0 0.52 0.78 0.53 0.28 0.31
OpenMVS [4] 0.13 0.77 0.12 0.80 0.11 0.82 0.13 0.80 0.65 0.56 1.02 0.53 0.74 0.58 0.28 0.31
@@
StreetSurf [11] 0.26 0.42 0.37 0.53 0.25 0.35 0.30 0.23 0.93 0.27 0.80 0.47 0.70 0.50 0.67 0.28
SCILLA (ours) 0.23 0.43 0.25 0.48 0.20 0.47 0.24 0.30 0.46 0.20 0.25 0.59 0.44 0.40 0.25 0.54
COLMAP [30] 0.27 0.56 0.78 0.53 0.19 0.67 0.27 0.58 0.67 0.48 0.85 0.67 0.80 0.62 0.80 0.68
OpenMVS [4] 0.3 0.57 0.76 0.53 0.21 0.67 0.24 0.63 0.62 0.51 0.86 0.69 0.76 0.70 0.75 0.69
@@ 4 Experiments
We evaluate the performance of SCILLA on multiple driving scenes and provide qualitative and quantitative comparisons compared to relevant methods.
4.1 Implementation details
@@
4.2 Datasets
We evaluate our method on four public driving datasets: KITTI-360 [20], nuScenes [3], Waymo Open Dataset [32] and Pandaset [41]. We conduct our experiments on four scenes from each dataset. Sequences were selected to have large differences between them. All sequences are static, except for Pandaset where dynamic vehicles are masked during training. Complete details about the selected sequences can be found in our supplementary.
4.3 Baseline
StreetSurf [11] is the current SoTA for surface reconstruction in driving scenes, using road-based surface initialization tailored for such sequences. It employs hierarchical space partitioning with 3 distinct fields to model close-range, far-range, and sky areas, along with 3D and 4D hash grid positional encoding and occupancy grids for efficient ray sampling. We used the official open-source implementation, evaluated with monocular depth supervision, and deactivated trajectory optimization for a fair comparison. All experiments were conducted on the same computational setup. We also extensively tested Neuralangelo [19] on the selected driving sequences with various training strategies and hyperparameters tuning but did not achieve satisfactory results. Neuralangelo’s need for a bounded scene to initialize the SDF in a sphere shape makes it unsuitable for the open, unbounded nature of driving sequences. Further details can be found in our supplementary materials. For a comprehensive comparison, we benchmark SCILLA’s mesh reconstruction against two classic MVS methods: Delaunay triangulation in COLMAP [30] and mesh reconstructionn [15] from OpenMVS [4], both using the denser point cloud from COLMAP.
4.4 Evaluation metrics
@@ 4.5 Results
Efficiency. We report in Tab. 2 the computational performance of each method. SCILLA uses a single 3D hierarchical hash grid, unlike StreetSurf’s two grids (3D and 4D) [11], making it more efficient with one-third of the parameters and half the GPU memory during training. Additionally, SCILLA requires half the training time per scene compared to StreetSurf.
@@
Table 2: Training analysis.
@@
The quantitative evaluation presented in the previous section (Sec. 4.5) is computed from the accumulated ground-truth LiDAR which is sparse and may miss some scene regions. To complete our evaluation, we present qualitative results in Fig. 8. Results demonstrate that SCILLA reconstructs higher-quality surfaces compared to StreetSurf and MVS methods. MVS-generated meshes have noisy normals and are incomplete. We observe that our method recovers more accurate and realistic scene details in many regions such as buildings, cars, poles, and trees. Our analysis supports the hypothesis that SteertSurf’s disentanglement of the close and far range based on the assumption that driving scenes are long and narrow is not adapted to the majority of driving sequences. In addition, StreetSurf’s road-surface initialization fails in sequences where the road and sidewalks are not at the same height or at non-flat roads such as downhills (see Fig. 8 – Seq. 30).
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
RGB (ref. only) COLMAP [30] OpenMVS [4] StreetSurf [11] SCILLA (ours)
@@
Figure 5: Qualitative experiments results on 8(a) KITTI-360s [20], 8(b) Pandaset [41], 8(c) nuScenes [3] and 8(d) Waymo Open Dataset [32]. RGB ground-truth images shown for reference of the scene layout. Complex scene geometries reconstructed in the mesh are highlighted in red squares. We compare our generated SDF mesh to COLMAP, OpenMVS and StreetSurf [11] meshes.
@@ 4.6 Ablation study
Convergence of density field compared to SDF field. We report in Fig. 6(a) the Chamfer Distance between LiDAR and predicted point clouds, for Seq. 31 from KITTI-360 and Seq. 916 from nuScenes, comparing two different NeRF: one solely based on density volumetric field and another one solely based on SDF at various training steps (without normal supervision, as we cannot supervise normal in case of pure density volumetric field). The results support the claim that density field converges faster compared to the SDF field.
SCILLA’s components. In this section, we ablate SCILLA’s SDF progressive initialization strategy, with quantitative results reported in Tab. 3. The results confirm that transitioning from volumetric to surface representation is essential for accurately initializing the SDF field and efficiently reconstructing driving sequences. We also report results with standard accumulated normal supervision and without network regularization loss.
@@
Table 3: Ablations study: we deactivate some SCILLA ’s key components. (a) without progressive initialization (b) with the accumulated normal for $\mathcal{L}_{\hat{N}}$ (c) without $\mathcal{L}_{\text{s}}$.
w/o SDF init w/o $\mathcal{L}_{\hat{N}}$ w/o $\mathcal{L}_{\text{s}}$ SCILLA
@@ 5 Discussion
Applications. Thanks to the quality and completeness of SCILLA’s reconstructed meshes, we can use modern MVS tools such as OpenMVS [4] to obtain a detailed and colorized representation of large urban scenes, as shown in Fig. 1. In addition, SCILLA is not restricted to driving scenarios, it can be applied to other types of scenes. We encourage the reader to refer to the supplementary materials for qualitative results of SCILLA on the Tanks&Temples dataset [18].
Limitations. Although our method can achieve highly detailed surface reconstruction in driving scenarios, it fails to reconstruct fine details in very wide and open sequences. We recommend consulting the supplementary materials for the failure cases of SCILLA.
@@ 6 Conclusion
In this work we presented SCILLA, an implicit hybrid approach to efficiently reconstruct large urban scenes. Our method’s hybrid architecture models the volumetric density and signed distance in two separate fields. We introduced a novel hybrid volume rendering strategy to progressively transition from volumetric representation to a signed distance representation. The quantitative and qualitative evaluations that we performed show that SCILLA achieves more precise reconstruction than state-of-the-art methods while being significantly faster and is better suited for a large variation of driving scenarios.
@@
In this supplementary material, we provide additional implementation details, experiments with Neuralangelo [19], additional qualitative results, ablation study on our samples attribution strategy, applications, and failure cases of SCILLA.
@@ 7 SDF-gradient normalization
As introduced in Sec. 3.3 of the main paper, volumetric representations based on density estimation tend to quickly approximate saturated alpha values ($\alpha_{i}^{v}$ being either 00 or $1$), while alpha values computed from the SDF converge slower. Because $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are composed jointly during the hybrid stage, we noticed that surface representation compensates for this gap by predicting large SDF gradient values so that $\alpha_{i}^{f}$ aligns with $\alpha_{i}^{v}$. Indeed, from Eq. 3 of the main paper, a solution to saturate alpha towards either 1 or 0, is to predict $f(p_{i+1})\ll f(p_{i})$ or $f(p_{i+1})\gg f(p_{i})$, respectively. During the forward pass, $f(p_{i})$ and $f(p_{i+1})$ are not directly predicted by the model but rather computed using:
@@
Neuralangelo [19] SCILLA (ours)
@@
Neuralangelo [19] SCILLA (ours)
@@
Figure 7: Qualitative experiments results on Seq. 30 7(a) and Seq. 31 7(b) from KITTI-360s [20]. Failure on Seq. 10061 7(c) Waymo Open Dataset [32]. We compare our generated SDF mesh to Neuralangelo [19] mesh for KITTI-360. We report the experiments on the full and segmented sequence for Waymo.

### v4

@@ 3.3 ViiNeuS’s volume rendering
This resulting color is then compared to the ground truth value to train our model. The strategy of associating a sample to either a volumetric or surface-based representation is discussed in the next section.
@@
SDF progressive initialization: Volumetric to surface representation.
Starting from a pure volumetric density training ($\alpha_{i}=\alpha_{i}^{v}$), we introduce a progressive initialization of the signed distance field that transitions from a volumetric to an SDF representation. Our approach is based on the assumption that estimating the density $\sigma_{i}$ at a specific location requires less complex geometric understanding, in contrast to estimating the signed distance to the closest point to a complex urban outdoor surface (i.e. $f(x_{i})$). The 3D geometry represented as a density field contains less information than the SDF field; the former contains occupancy information (which can be considered as binary) whereas the latter represents, additionally, continuous distances to the surface, including close and far-range points as well as the empty spaces between driving scene objects, such as cars and trees. Therefore, in this specific scenario where observations have reduced overlap introducing ambiguity in the geometry estimation, learning a density field is simpler compared to the SDF field (see figure 5 of our ablation section). Furthermore, by initially leveraging a pure volumetric density representation during training, we shorten the convergence time of our density estimator.
@@
  • Volumetric stage: All samples along the ray express volumetric densities, and their alpha values are computed with $\sigma_{i}$.
  • Hybrid stage: we gradually assign samples along the ray to train our signed distance field. Within the same ray, a portion of the sample values is represented by volumetric alpha $\alpha_{i}^{v}$, while the remainder is computed with $\alpha_{i}^{f}$.
@@
Hybrid stage regularization.
Composing alpha values obtained from our two different geometric representations is not straightforward as the learning dynamic of the density and the SDF are not the same. As mentioned earlier and due to the partial observability of the reconstructed environment, learning a volumetric representation of the scene based only on spatial density is easier and faster compared to learning an arbitrary signed distance field of a complex urban scene. To facilitate the convergence of our representation, we introduce two regularization methods: probability-based sampling attribution and SDF-gradient normalization.
Probability-based samples attribution. To assign the alpha value based on density or SDF to a sample, we initially start with a random selection for each ray. We noticed that this association strategy is not suitable: the signed distance field could not be properly initialized during the hybrid stage. To solve this problem, we introduce a probability-based sample association. We assign $\alpha_{i}^{f}$ to samples with the highest predicted density to initialize the signed distance field. It facilitates the learning of the signed distance field by limiting the range of predicted values, as all the points are sampled close to a surface. By sampling points near the surface, we handle the wide range of distances of driving scenarios and complex geometries.
SDF-gradient normalization. During the hybrid stage, where $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are jointly composed, we observed that the surface representation often predicts large SDF gradient values to align $\alpha_{i}^{f}$ with $\alpha_{i}^{v}$ (for more details, see our supplementary materials). To address this, we normalize the SDF gradient in the $\alpha_{i}^{f}$ computation to prevent the gradient from compensating the alpha distribution difference.
@@ 3.4 Optimization
According to NeuS formulation of the alpha compositing equation (Eq. 3), the network converges properly when the standard deviation $1/s$ of $\phi_{s}$ approaches zero. Due to the complexity of driving scenarios and the large scale of urban scenes, the parameter $s$ doesn’t always reach a satisfactory level of convergence. Unlike StreetSurf which manually fixes the parameter $s$ to increment linearly, we introduce a regularization term on $s$ during training: $\mathcal{L}_{\text{s}}=1/(s+\epsilon)$.
@@
StreetSurf [12] 0.14 0.50 0.09 0.71 0.10 0.67 0.11 0.66 2.33 0.10 0.47 0.40 0.36 0.30 0.27 0.31
ViiNeuS (ours) 0.13 0.56 0.11 0.71 0.11 0.66 0.13 0.72 0.49 0.23 0.28 0.46 0.33 0.29 0.35 0.34
COLMAP [32] 0.11 0.70 0.09 0.81 0.08 0.78 0.08 0.74 0.53 0.41 0.85 0.52 0.65 0.53 0.26 0.63
OpenMVS [26] 0.11 0.77 0.08 0.80 0.08 0.82 0.10 0.80 0.62 0.56 1.0 0.53 0.73 0.58 0.31 0.64
GOF [47] - sparse 0.29 0.53 0.23 0.63 0.45 0.51 0.90 0.54 0.42 0.59 0.51 0.45
GOF [47] - dense 0.17 0.71 0.16 0.72 0.20 0.74 0.11 0.80 0.51 0.48 0.86 0.62 0.41 0.67 0.30 0.55
@@
StreetSurf [12] 0.22 0.43 0.35 0.53 0.23 0.35 0.25 0.24 0.57 0.29 0.78 0.47 0.67 0.50 0.65 0.28
ViiNeuS (ours) 0.19 0.44 0.22 0.48 0.14 0.47 0.19 0.30 0.40 0.20 0.22 0.59 0.40 0.40 0.22 0.54
COLMAP [32] 0.18 0.57 0.71 0.53 0.14 0.67 0.19 0.59 0.25 0.51 0.77 0.67 0.65 0.62 0.70 0.68
OpenMVS [26] 0.25 0.58 0.76 0.53 0.20 0.67 0.22 0.64 0.28 0.54 0.85 0.69 0.75 0.70 0.74 0.69
GOF [47] - sparse 1.87 0.32 2.32 0.20 1.63 0.36 1.54 0.29 1.55 0.07 1.72 0.16 1.49 0.12 1.41 0.18
GOF [47] - dense 1.20 0.38 1.17 0.39 1.55 0.41 2.11 0.34 1.02 0.12 1.55 0.23 1.44 0.12 1.06 0.29
@@ 4 Experiments
We evaluate the performance of ViiNeuS on multiple driving scenes and provide qualitative and quantitative comparisons compared to relevant methods.
4.1 Implementation details.
@@
4.2 Datasets.
We evaluate our method on four public autonomous driving datasets: KITTI-360 [21], nuScenes [3], Waymo Open Dataset [34] and Pandaset [43]. We conduct our experiments on four static scenes from each one of these datasets. Since our method is designed for reconstructing the static background of urban areas with limited image overlap, we mask dynamic vehicles and pedestrians in the Pandaset sequences during training. Further details on the selected sequences are provided in the supplementary materials. We use poses provided by the datasets except for Waymo where we recompute the vehicle trajectory and sensor calibration with MOISST [14] as the ones provided are not accurate.
4.3 Baseline.
We evaluate the state-of-the-art (SoTA) method StreetSurf [12] for driving scenes surface reconstruction, using its official implementation with monocular depth supervision and trajectory optimization disabled for a fair comparaison. All experiments were conducted on the same computational setup. GOF [47], a Gaussian splitting method for landmark surface reconstruction, was also evaluated using the official implementation. We modified it to initialize the Gaussians with a dense point cloud from COLMAP and adjusted the rasterizer for cameras with an off-center focal point. We evaluate both sparse and dense point cloud initialization. We extensively tested Neuralangelo [20] on the selected driving sequences, with various training strategies and hyperparameters tuning. However we did not obtain satisfactory results: Neuralangelo requires a bounded scene to initialize the SDF in a sphere shape, making it unsuitable for the open, complex nature of driving sequences. For more details on these experiments, we refer readers to our supplementary materials. To provide a more comprehensive comparison we benchmark ViiNeuS mesh reconstruction with meshes obtained by two methods based on classic MVS: Delaunay triangulation implemented in COLMAP [32] and mesh reconstruction [17] from OpenMVS [26]. For both of the MVS meshing baselines, we used the dense point cloud generated by COLMAP as it provides denser results compared to the one obtained with OpenMVS library.
4.4 Evaluation metrics.
@@ 4.5 Results
[…]
Efficiency. We report in Tab. 2 the computational performance of each method. ViiNeuS is more efficient than StreetSurf [12], using a single 3D hierarchical hash grid instead of two grids (3D and 4D), with a third of the parameters and half the GPU memory. GOF, on the other hand, has variable GPU memory and parameters depending on the number of Gaussians per scene. For a single scene, ViiNeuS requires half the training time compared to StreetSurf (20 min vs 40 min), while GOF takes at least 1.5 hours, including 30 minutes each for point cloud initialization, training, and mesh extraction.
@@
Table 2: Training analysis: reported using the same GPU, equivalent to an RTX4090. Training time excludes COLMAP point cloud initialization.
@@
The quantitative evaluation presented in the previous section (Sec. 4.5) is computed from the accumulated ground-truth LiDAR which are sparse and do not cover some regions of the scene. To complete our evaluation, we present qualitative results in Figs.  8. Results demonstrate that ViiNeuS reconstructs higher-quality surfaces compared to StreetSurf and explicit MVS methods. Even if finer details can be recovered on some parts of the scene, meshes obtained by explicit MVS methods have noisy normals and are incomplete, which limits their applicability to some tasks such as mesh texturing (see supplementary materials). We observe that our method recovers more accurate and realistic scene details at many regions such as buildings, cars, poles, and trees. Our analysis supports the hypothesis that SteertSurf’s disentanglement of the close and far range based on the assumption that driving scenes are long and narrow is not adapted to the majority of driving sequences. In addition, StreetSurf’s road-surface initialization fails in sequences where the road and sidewalks are not at the same height or at non-flat roads such as downhills (see Fig. 8 – Seq. 30).
COLMAP [32] OpenMVS [26] GOF [47] StreetSurf [12] ViiNeuS (ours)
@@
COLMAP [32] OpenMVS [26] GOF [47] StreetSurf [12] ViiNeuS (ours)
@@
Figure 4: Qualitative experiments results on 8(a) KITTI-360, 8(b) Pandaset, 8(c) nuScenes and 8(d) Waymo Open Dataset. We compare our mesh extracted from our SDF to GOF, COLMAP, OpenMVS and StreetSurf meshes.
@@ 4.6 Ablation study
Convergence of density field compared to SDF field.
We report in Fig. 5 the Chamfer Distance between LiDAR and predicted point clouds, for Seq. 31 from KITTI-360 and Seq. 916 from nuScenes, comparing two different NeRF: one solely based on density volumetric field and another one solely based on SDF at various training steps (without normal supervision, as we cannot supervise normal in case of pure density volumetric field). The results support the claim that density field converges faster compared to the SDF field.
Duration of hybrid stage.
@@
ViiNeuS’s components.
We ablate the effect of ViiNeuS’s SDF progressive initialization strategy and show quantitative results in Tab. 3. The results demonstrate that our initialization that transitions from the volumetric to the surface representation is crucial to accurately initialize the SDF field and efficiently reconstruct driving sequences. In addition, we report our method with standard accumulated normal supervision, and without the network regularization loss.
Table 3: Ablations study: we deactivate some of the key components of ViiNeuS in order to measure the impact on the final reconstruction. (a) without progressive initialization (b) with the accumulated normal for $\mathcal{L}_{\hat{N}}$ (c) without $\mathcal{L}_{\text{s}}$.
w/o SDF init w/o $\mathcal{L}_{\hat{N}}$ w/o $\mathcal{L}_{\text{s}}$ ViiNeuS
@@ 5 Discussion
Applications. Thanks to the quality and completeness of ViiNeuS’s reconstructed meshes, we can use modern MVS tools such as OpenMVS [26] to obtain a detailed and colorized representation of large urban scenes, as shown in Fig. 1.
Limitations. Although our method can achieve highly detailed surface reconstruction in driving scenarios, it fails to reconstruct fine details in very wide and open sequences. We recommend consulting the supplementary materials for the failure cases of ViiNeuS.
@@ 6 Conclusion
In this work we presented ViiNeuS, an implicit hybrid approach to efficiently reconstruct large urban scenes. Our method’s hybrid architecture models the volumetric density and signed distance in two separate fields. We introduced a novel hybrid volume rendering strategy to progressively transition from volumetric representation to a signed distance representation. The quantitative and qualitative evaluations that we performed show that ViiNeuS achieves more precise reconstruction than state-of-the-art methods while being significantly faster and is better suited for a large variation of driving scenarios.
@@
In this supplementary material, we provide additional implementation details, experiments with Neuralangelo [20] and additional quantitative and qualitative results. Furthermore, an ablation study on our samples attribution strategy, applications, and failure cases of ViiNeuS are also provided.
@@ 7 SDF-gradient normalization
As explained in Sec. 3.3 of the main paper, volumetric representations based on density estimation tend to quickly approximate saturated alpha values ($\alpha_{i}^{v}$ being either 00 or $1$), while alpha values computed from the SDF converge slower. Because $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are composed jointly during the hybrid stage, we noticed that the surface representation compensates for this gap by predicting large SDF gradients so that $\alpha_{i}^{f}$ aligns with $\alpha_{i}^{v}$. Indeed, from Eq. 3 of the main paper, a solution to saturate alpha towards either 1 or 0, is to predict $f(p_{i+1})\ll f(p_{i})$ or $f(p_{i+1})\gg f(p_{i})$, respectively. During the forward pass, $f(p_{i})$ and $f(p_{i+1})$ are not directly predicted by the model but rather computed using:
@@
Neuralangelo [20] ViiNeuS (ours)
@@
Neuralangelo [20] ViiNeuS (ours)
@@
Figure 6: Qualitative experiments results on Seq. 30 6(a) and Seq. 31 6(b) from KITTI-360s [21]. Failure on Seq. 10061 6(c) Waymo Open Dataset [34]. We compare our generated SDF mesh to Neuralangelo [20] mesh for KITTI-360. We report the experiments on the full and segmented sequence for Waymo.